OpenStudy (anonymous):

A population of bacteria is growing according to the function below, where t is in hours. How many hours will it take for the population to grow to 1,000,000 bacteria? B(t) = 0.5 e^1.5t

OpenStudy (anonymous):

Well first we set up this equation like this: 1000000=0.5e^(1.5t)

OpenStudy (anonymous):

now we have to solve for t

OpenStudy (anonymous):

I was able to get that far, but forgot where to go after that! Help?

OpenStudy (anonymous):

Now, what you want to do is divide both sides by 0.5

OpenStudy (anonymous):

OK - I get 2,000,000= e^1.5t now what?

OpenStudy (anonymous):

2000000=e^1.5t

OpenStudy (anonymous):

now, here is the trick we need to take the natural log of both sides

OpenStudy (anonymous):

remind me of that one please, is that e?

OpenStudy (anonymous):

lne^1.5t=ln(2000000)

OpenStudy (anonymous):

now , the ln and the e will cancel out, since they are inverses of each other, and you well get: 1.5t=ln(2000000)

OpenStudy (anonymous):

Now, simply divide both sides by 1/5 to isolate t: t=1n(2000000)/1.5

OpenStudy (anonymous):

OK so when I do that I get 9.672438492 How close am I?

OpenStudy (anonymous):

that is the answer, nice job.

OpenStudy (anonymous):

THANK YOU!!!!!!